The area of a rectangle is 54 cm2. If the length of the rectangle is 3 more than its width, find the length and width of the rectangle.
Answer by edjones(8007)

L*W=A
L=w+3
w(w+3)=54
w^2+3w=54
w^2+3w-54=0
Factor: (w+9)(w-6)=0
w=-9 w=6 w can't be -9 so thow -9 out.
w=6 cm
Check:
(6+3)*6=54 true
Ed